OpenStudy (anonymous):

Marta's starting annual salary is $24, 900 At the beginning of each new year, she receives a $2700 raise. Write a recursive formula to find Marta's salary f(n) after n years. What will Marta's salary be after 4 years? a) f(n) = 2700f(n - 1); $33, 000 b) f(n) = f(n - 1) + 2700; $33, 000 c) f(n) = f(n - 1) - 2700; $35, 700 d) f(n) = -2700f(n - 1); $35, 700

OpenStudy (ranga):

$2700 raise each year. Salary any year = previous year's salary + $2,700 f(n) = f(n-1) + 2,700

OpenStudy (ranga):

Salary after 1 year: 24,900 + 2700 = 27,600 after 2 years: 27,600 + 2700 = 30300 after 3 years: 30300 + 2700 = 33000

OpenStudy (ranga):

But after 4 years it should be 33000 + 2700 = 35700

OpenStudy (ranga):

b) is the correct answer for the recursive equation. But the salary after 4 years comes to $35,700 and not $33,000!!!

OpenStudy (ranga):

Oh, they are counting n = 0 as the first year. So $33,000 is fine. So answer b). Go ahead post them.

OpenStudy (anonymous):

The function b(n) =6n represents the number of light bulbs b(n) that are needed for n chandeliers. How many light bulbs are needed for 15 chandeliers? Choose one answer. a. 90 light bulbs b. 2 light bulbs c. 96 light bulbs d. 80 light bulbs

OpenStudy (ranga):

b(n) =6n where n is number of chandeliers. To find the number of light bulbs needed for 15 chandeliers, put n = 15: b(15) = 6 * 15 = 90 light bulbs

OpenStudy (anonymous):

and this is the one i need you to check. Suppose your business has a special checking account used just for paying the phone bill. The balance is $740.00 this month. Next month the balance will be $707.60, after that it will be $675.20, and on the third month the balance will be $642.80. Write an explicit formula to represent the balance in the account as an arithmetic sequence. How many months can you pay your phone bill without depositing any more money in the account? a) A(n) = 740.00 - 32.40n; 22 months. b) A(n) = 740.00 + (n - 1)(-32.40); 23 months. c) A(n) = 740.00 - 32.40n; 23 months. d) A(n) = 740.00 + (n - 1)(-32.40); 24 months. Is B the answer

OpenStudy (anonymous):

So the one answer is A, 90 light bulbs?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

Okay. I posted the one question. Is the answer B to the last question i posted?

OpenStudy (ranga):

Yes b) is the correct answer for the last one.

OpenStudy (anonymous):

Thank you! So kid answered this question for me and I was wondering if it was also right. Tell whether the sequence is arithmetic. If it is, what is the common difference? 15, 19, 23, 27, . . . a. yes; 15/19 b. yes; 4 c. yes; 34 d. no He choose b.

OpenStudy (anonymous):

Some*

OpenStudy (ranga):

Yes. It is an arithmetic sequence with common difference 4.

OpenStudy (anonymous):

Are these correct also. I hate asking you. but, it's my last quiz and I want to get a good grade. So i am making sure. 1. Describe a pattern in each sequence. What are the next two terms of each sequence? 24, 22, 20, 18, . . . Choose one answer. a. multiply the previous term by -2; -36, 72 b. multiply the previous term by -2; 16, 72 c. subtract 2 from the previous term; 16, 14 d. add 2 to the previous term; 20, 22 I choose C. 2. Describe a pattern in each sequence. What are the next two terms of each sequence? 2, 4, 8, 16, . . . Choose one answer. a. add 2 to the previous term; 14, 12 b. subtract 2 from the previous term; 32, 64 c. multiply the previous term by -2; -32, 64 d. multiply the previous term by 2; 32, 64 i choose D.

OpenStudy (ranga):

both are correct.

OpenStudy (anonymous):

Okay thank you so much for your help, I know who to come to if i ever need help.

OpenStudy (ranga):

Alright!!! You are welcome.